OpenStudy (anonymous):

As soon as a traffic light turns green, a car speeds up from rest to 58.0 mi/h with constant acceleration 10.00 mi/h/s. In the adjoining bike lane, a cyclist speeds up from rest to 19.0 mi/h with constant acceleration 14.50 mi/h/s. Each vehicle maintains constant velocity after reaching its cruising speed. (a) For what time interval is the bicycle ahead of the car? s (b) By what maximum distance does the bicycle lead the car? ft

OpenStudy (dumbcow):

First convert the rates to ft/s instead of mi/hr, since we want to evaluate in terms of seconds 1 mi/hr = 22/15 ft/s Also note that a(t) = d/dt v(t) and v(t) = d/dt x(t) , a is acceleration v is velocity, and x is position Let t* represent time it takes to reach cruising speed Bike: a = 21.26 , v = 21.26t, x = 10.63t^2 ; when t <t* v(t*) = 27.86 = 21.26t* --> t* = 1.31 It take the bicyclist 1.31 sec to reach cruising speed of 27.86 ft/s x(t*) = 10.63(1.31^2) = 18.26 After 1.31 sec the cyclist has traveled 18.26 ft x = 18.26 +27.86(t - 1.31) ; when t > t* Car: a = 14.66 , v = 14.66t, x = 7.33t^2 ; when t <t* v(t*) = 85.07 = 14.66t* --> t* = 5.8 It take the car 5.8 sec to reach cruising speed of 85.07 ft/s x(t*) = 7.33(5.8^2) = 246.8 After 5.8 sec the car will have traveled 246.8 ft x = 246.8 +85.07(t - 5.8) ; when t > t* Now when t < 1.31, we know that the bicycle is ahead of the car --> 10.63t^2 > 7.33t^2 At t=1.31 the cyclist stops accelerating and then by setting their positions equal we can find the time t that the car catches up with the cyclist --> 7.33t^2 = 18.26+27.86(t - 1.31) solve the quadratic 7.33t^2 -27.86t +18.26 = 0 using quadratic formula t = 0.84 or 2.96, however we know that t must be greater than 1.31 t = 2.96 part a) The cyclist is ahead of the car for first 2.96 seconds For next part to find max distance, define distance d d = x_cyclist - x_car d = 18.26+27.86(t - 1.31) - 7.33t^2 d = -7.33t^2 +27.86t -18.26 Now find vertex of the parabola or find max by setting derivative equal to 0. dd/dt = -14.66t +27.86 = 0 t = 1.9 --> max_d = 8.21 feet